5. State the dolgain, range, x-intercept(s) y-intercept(s) and asymptote(s) of the given function. (7 marks) y=9(3*) - 1
Answer #1


y = 9(3%). at a intercepts y=0 So 0 = 90334-1 37 3 11 1 हा 3 32 a -2 x intercepts is (-2,0) at y interrepte 0 -) Y = 9(3)° )